Warriorz13:

A 50.0 g sample of 16N decays to 12.5g in 14.4 seconds. What is its half-life?

Warriorz13:

@JustSaiyan

JustSaiyan:

50.0g ÷ 2 = 25 = 1 half life 25 ÷ 2 = 12.5g = 2 half-lives; 14.4 ÷ 2 = 7.2 sec = half-life

Warriorz13:

So what is the half life? 7.2 seconds?

JustSaiyan:

Yes.